Good Day Online Tutors. Can you please help me about this problem ? Thank you☺
Kageyama has 3,000 pesos consisting of 20 pesos, 50 pesos, 100 pesos and 200 peso bills. The number of 20 peso bills is the same as the number of 100 peso bills. The number of 200 peso bills is twice the number of 100 peso bills. The number of 50 peso bills is three more than the number of 20 peso bills. How many each type of bill does he have?

Answer by ikleyn(53763)

.
Kageyama has 3,000 pesos consisting of 20 pesos, 50 pesos, 100 pesos and 200 peso bills.
The number of 20 peso bills is the same as the number of 100 peso bills.
The number of 200 peso bills is twice the number of 100 peso bills.
The number of 50 peso bills is three more than the number of 20 peso bills.
How many each type of bill does he have?
~~~~~~~~~~~~~~~~~~

Let  x   be the number of 20-pesos bills

Then from the problem you have

    x = the number of the 100-pesos bills

    2x = the number of the 200-pesos bills

    (x+3) = the number of the 50-pesos bills



Now you write the total money equation

    20x + 50*(x+3) + 100x + 200*(2x) = 3000.



Simplify and solve


    20x + 50x + 100x + 400x = 3000-150

            570x            = 2850

               x            = 2850/570 = 5.



ANSWER.  5 20-pesos bills;  5 100-peses bills;  10  200-pesos bills;  8 50-pesos bills.

Solved (with using only one unknown and one equation)